# Continuity of the continued fraction mapping revisited

### v1

@@ 1. Introduction
Given $x\in[0,1)$, we define
(1.1) $d_{1}(x)\coloneqq\begin{cases}\lfloor 1/x\rfloor,&\text{if }x\neq 0;\\ \infty,&\text{if }x=0,\end{cases}\quad\text{ and }\quad T(x)\coloneqq\begin{cases}1/x-\lfloor 1/x\rfloor,&\text{if }x\neq 0;\\ 0,&\text{if }x=0,\end{cases}$
where $\lfloor\xi\rfloor$ denotes the largest integer not exceeding $\xi\in\mathbb{R}$, and define
(1.2) $d_{n}(x)\coloneqq d_{1}(T^{n-1}(x))$
for each $n\in\mathbb{N}$. The $d_{n}(x)$ are called the partial quotients of the continued fraction expansions. It is well-known that the algorithms (1.1) and (1.2) yield a unique continued fraction expansion
$x=[d_{1}(x),d_{2}(x),\dotsc]\coloneqq\cfrac{1}{d_{1}(x)+\cfrac{1}{d_{2}(x)+\ddots}},$
with the conventions $1/\infty=0$ and $\infty+0=\infty$. For irrational $x\in[0,1)$, we have $(d_{k}(x))_{k\in\mathbb{N}}\in\mathbb{N}^{\mathbb{N}}$, while for rational $x\in(0,1)$, we have $n\coloneqq\inf\{k\in\mathbb{N}:d_{k}(x)=\infty\}<\infty$ so that $x=[\underbrace{d_{1}(x),\dotsc,d_{n-1}(x)}_{\text{$n-1$ terms}},\infty,\infty,\dotsc]$ with $(d_{k}(x))_{k=1}^{n-1}\in\mathbb{N}^{n-1}$ (see [5, p. 4]).
The mapping from $[0,1)$ to $(\mathbb{N}\cup\{\infty\})^{\mathbb{N}}$ defined by
$x\mapsto(d_{1}(x),d_{2}(x),d_{3}(x),\dotsc)$
for each $x\in[0,1)$ is referred to as the continued fraction mapping. A classical result in the study of continued fractions is that the set of irrationals in $[0,1)$, equipped with the subspace topology inherited from the Euclidean space $\mathbb{R}$, and the product space $\mathbb{N}^{\mathbb{N}}$, where $\mathbb{N}$ is a discrete space, are homeomorphic via the continued fraction mapping.
In this paper, building upon our previous work [1, Section 3.2], where we studied the continuity of the Pierce expansion mapping, we address the continuity problem of the continued fraction mapping on $[0,1)$, encompassing rational numbers, and adjust the codomain and its topology accordingly. To achieve this, in Section 2, we provide some basic results on continued fractions. In Section 3, as the main results of this paper, we establish the continuity result for an extension of the continued fraction mapping to $[0,1]$ (see Theorem 3.12) and the continuity result for the original continued fraction mapping on $[0,1)$ (see Corollary 3.13).
@@ 2. Preliminaries of continued fractions
In this section, we list some basic properties of continued fractions. Most of the results in this section can be found in the classical textbook on number theory by Hardy and Wright [4], as well as in books more focused on the theory of continued fractions by Iosifescu and Kraaikamp [5] and Khinchin [Khi97].
We slightly extend the domain of the continued fraction mapping to the closed unit interval $[0,1]$. For each $x\in[0,1]$, define
@@
By definitions, $\overline{T}=T$ and $\overline{d}_{n}=d_{n}$ for every $n\in\mathbb{N}$, on the original domain $[0,1)$. Define $f\colon[0,1]\to\mathbb{N}_{\infty}^{\mathbb{N}}$ by
@@
for each $x\in[0,1]$. Clearly, $f$ is well-defined by definition, and $f$ is an extension of the continued fraction mapping in the sense that $f(x)=(d_{k}(x))_{k\in\mathbb{N}}$ for any $x\in[0,1)$. We call $f$ the extended continued fraction mapping.
@@
$\widetilde{\varphi}_{k}(\sigma)\coloneqq[\underbrace{\sigma_{1},\sigma_{2},\dotsc,\sigma_{k}}_{\text{$k$ terms}},\infty,\infty,\dotsc]$
for each $k\in\mathbb{N}_{0}$. Now, define $\widetilde{\varphi}\colon\mathbb{N}_{\infty}^{\mathbb{N}}\to[0,1]$ by
@@
$\varphi_{n}(\sigma)\coloneqq\frac{p_{n}^{*}(\sigma)}{q_{n}^{*}(\sigma)},$
@@
We remark that if the algorithms (1.1) and (1.2) were absent, then every rational number in $(0,1)$ would admit two distinct continued fraction expansions (see [4, Theorem 162] or [5, p. 3]). This is because, for any integer $k\geq 2$, we have $1/k=1/[(k-1)+1/1]$ with $k-1\in\mathbb{N}$. The following proposition rephrases this fact in terms of two maps $f\colon[0,1]\to\Sigma$ and $\varphi\colon\Sigma\to[0,1]$.
@@
    $\sigma$ $\coloneqq(\underbrace{\overline{d}_{1}(x),\overline{d}_{2}(x),\dotsc,\overline{d}_{n-1}(x)}_{\text{$n-1$ terms}},\overline{d}_{n}(x))=f(x)\in\Sigma_{n}\cap f([0,1]),$
    $\sigma^{\prime}$ $\coloneqq(\underbrace{\overline{d}_{1}(x),\overline{d}_{2}(x),\dotsc,\overline{d}_{n-1}(x)}_{\text{$n-1$ terms}},\overline{d}_{n}(x)-1,1)\in\Sigma_{n+1}\setminus f([0,1]),$
@@
We further define the fundamental interval associated with $\sigma$ by
$I_{\sigma}\coloneqq f^{-1}(\Upsilon_{\sigma})=\{x\in[0,1]:\overline{d}_{k}(x)=\sigma_{k}\text{ for all }k\in\{1,\dotsc,n\}\}.$
It is legitimate to call $I_{\sigma}$ as an interval due to the following proposition.
Proposition 2.5 (See [5, Theorem 1.2.2]).
For each $n\in\mathbb{N}$ and $\sigma\coloneqq(\sigma_{k})_{k\in\mathbb{N}}\in\Sigma_{n}$, the set $I_{\sigma}$ is an interval with endpoints $\varphi(\sigma)$ and $\varphi(\widehat{\sigma})$, where $\widehat{\sigma}\in\Sigma_{n}$ is given by
$\widehat{\sigma}\coloneqq(\underbrace{\sigma_{1},\dotsc,\sigma_{n-1}}_{\text{$n-1$ terms}},\sigma_{n}+1,\infty,\infty,\dotsc).$
@@ 3. Main results and proofs
For the first assertion, we just note that $\sum_{k=1}^{\infty}1/F_{k}^{2}<\infty$, which is clear from Binet’s formula (1.3). We omit the remaining details, as the proof follows the same line as the standard proof that any countable product of metric spaces is metrizable. The second statement follows from Tychonoff’s theorem, which tells us that the product space $\mathbb{N}_{\infty}^{\mathbb{N}}$ is compact. ∎
@@
Note first that, for any $\sigma\in\mathbb{N}_{\infty}^{\mathbb{N}}$, we have $g(\sigma)\in\Sigma$, and so $\varphi(g(\sigma))=\widetilde{\varphi}(g(\sigma))$ by definition of $\varphi$ in (2.4). If $\sigma\in\Sigma_{\infty}$, then $g(\sigma)=\sigma$ by definition, and hence, $\widetilde{\varphi}(g(\sigma))=\widetilde{\varphi}(\sigma)$. Now, assume $\sigma\in\mathbb{N}_{\infty}^{\mathbb{N}}\setminus\Sigma_{\infty}$, and put $\sigma\coloneqq(\sigma_{k})_{k\in\mathbb{N}}$. Then, we can find the unique $n\in\mathbb{N}_{0}$ such that $\sigma^{(n)}\in\Sigma_{n}$ and $\sigma_{n+1}=\infty$, so that $g(\sigma)=\sigma^{(n)}\in\Sigma_{n}$. It follows by definition that
$\widetilde{\varphi}(g(\sigma))=[\sigma_{1},\dotsc,\sigma_{n},\infty,\infty,\dotsc]=[\sigma_{1},\dotsc,\sigma_{n},\infty,\sigma_{n+2},\sigma_{n+3},\dotsc]=\widetilde{\varphi}(\sigma),$
and this completes the proof. ∎
@@
Let $\sim_{g}$ be a binary relation on $\mathbb{N}_{\infty}^{\mathbb{N}}$ given by $\sigma\sim_{g}\tau$ if and only if $g(\sigma)=g(\tau)$ for $\sigma,\tau\in\mathbb{N}_{\infty}^{\mathbb{N}}$. Evidently, $\sim_{g}$ is an equivalence relation on $\mathbb{N}_{\infty}^{\mathbb{N}}$. Note that the map $\widetilde{g}\colon\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}\to\Sigma$ given by $\widetilde{g}([\sigma])=g(\sigma)$, for each $[\sigma]\in\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}$, is a bijection. We equip $\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}$ with the quotient topology, and define
@@
Case I. Assume $\tau_{M}\neq\infty$. Since the first $M$ terms of $g(\sigma)$ are all finite, we have $\sigma^{(M)}=((g(\sigma))^{(M)}\in\Sigma_{M}$ by definition of $g$; similarly, $\tau^{(M)}=((g(\tau))^{(M)}\in\Sigma_{M}$. Note that $\sigma_{M}\neq\tau_{M}$ by definition of $M$. Since $\bm{\sigma}_{k}\to\sigma$ and $\bm{\tau}_{k}\to\tau$ as $k\to\infty$ in the product space $\mathbb{N}_{\infty}^{\mathbb{N}}$, we know that $(\bm{\sigma}_{k})^{(M)}=\sigma^{(M)}\in\Sigma_{M}$ and $(\bm{\tau}_{k})^{(M)}=\tau^{(M)}\in\Sigma_{M}$ for all large enough $k$. Then, for such $k$, it must be that $(g(\bm{\sigma}_{k}))_{M}=\sigma_{M}$ and $(g(\bm{\tau}_{k}))_{M}=\tau_{M}$. But $\sigma_{M}\neq\tau_{M}$, and it follows that $g(\bm{\sigma}_{k})\neq g(\bm{\tau}_{k})$, i.e., $\bm{\sigma}_{k}\not\sim_{g}\bm{\tau}_{k}$, for all large enough $k$, a contradiction.
@@
Let $X$ be a compact space and $\sim$ an equivalence relation on $X$. If $X/\sim$ denotes the quotient space, then the following are equivalent:
@@
We first note that by Lemma 3.7 and Proposition 3.8, the canonical projection $\pi_{g}\colon\mathbb{N}_{\infty}^{\mathbb{N}}\to\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}$ is a closed mapping. But $g=\widetilde{g}\circ\pi_{g}$ by Lemma 3.6, where $\widetilde{g}$ is a homeomorphism (Lemma 3.5), and this implies that $g$ is also a closed mapping.
@@
    $\lim_{\begin{subarray}{c}t\to x\\ t\in I_{\sigma}\end{subarray}}f(t)=\sigma\quad\text{and}\quad\lim_{\begin{subarray}{c}t\to x\\ t\not\in I_{\sigma}\end{subarray}}f(t)=\tau.$
@@
(i) See the proof of [1, Lemma 3.9]. The proof follows the same line, using the compact metrizability of $(\Sigma,\mathcal{T}_{\Sigma})$ (Lemma 3.10) and the fact that $\varphi^{-1}(\{x\})$ is a singleton for each $x\in\mathbb{I}\cup\{0,1\}$ (Proposition 2.4(i)).
(ii) See the proof of [1, Lemma 3.10]. The proof follows the same line, using the compact metrizability of the subspaces $\Upsilon_{\sigma}$ and $\Sigma\setminus\Upsilon_{\sigma}$ of $(\Sigma,\mathcal{T}_{\Sigma})$ (Lemma 3.11) and the fact that $\varphi^{-1}(\{x\})$ is a doubleton for each $x\in(0,1)\cap\mathbb{Q}$ (Proposition 2.4(ii)). ∎
Corollary 3.13.
The continued fraction mapping on $[0,1)$ to $(\Sigma,\mathcal{T}_{\Sigma})$ defined by $x\mapsto(d_{1}(x),d_{2}(x),\dotsc)$ for each $x\in[0,1)$ is:
  1. (i)
    Continuous at every irrational point and at 00.
  2. (ii)
    One-sided continuous only at every rational point in $(0,1)$; precisely, if we let $x\in(0,1)\cap\mathbb{Q}$ and put $\varphi^{-1}(\{x\})=\{\sigma,\tau\}$, then
    $\lim_{\begin{subarray}{c}t\to x\\ t\in I_{\sigma}\end{subarray}}f(t)=\sigma\quad\text{and}\quad\lim_{\begin{subarray}{c}t\to x\\ t\not\in I_{\sigma}\end{subarray}}f(t)=\tau.$
Proof.
The corollary is an immediate consequence of Theorem 3.12. ∎
Theorem 3.14.
Proof.
$\iff\varphi^{-1}(U)\text{ is open in }(\Sigma,\mathcal{T}_{\Sigma}),$
@@
It is a standard fact that the sets $\mathbb{I}$ and $\Sigma_{\infty}$ are in a bijective relation via the continued fraction mapping. Therefore, to obtain the corollary, we only need to combine Theorems 3.12(i) and 3.14. ∎

### v2

@@ 1. Introduction
Given $x\in(0,1)$, the continued fraction expansion is expressed as:
$x=[d_{1}(x),d_{2}(x),\dotsc]\coloneqq\cfrac{1}{d_{1}(x)+\cfrac{1}{d_{2}(x)+\cfrac{1}{d_{3}(x)+\ddots}}},$
where the $d_{n}(x)$ are the partial quotients. For irrationals, this expansion is infinite, forming sequences in $\mathbb{N}^{\mathbb{N}}$; for rational numbers, the expansion terminates after a finite number of terms, say $n$, with the condition $d_{n}(x)\geq 2$. This dual behavior motivates a detailed exploration of the continuity of the continued fraction mapping across the unit interval.
While the continued fraction mapping is a homeomorphism when restricted to the irrationals, its behavior at rational points is discontinuous in the classical sense. By redefining the codomain topology, one-sided continuity can be established at rational points. This paper develops such a topological framework, demonstrating how it simplifies analysis while maintaining rigor.
Topological insights into continued fractions have practical implications. For instance, Ridley and Petruska [7] introduced the error-sum function of continued fractions and studied it directly on the unit interval, requiring heavy machinery and complex calculations. In [1], we investigated the error-sum function in the context of Pierce expansions, leveraging the topological viewpoint to simplify the analysis. This approach significantly reduced computational complexity while retaining mathematical rigor. Similarly, in the context of continued fractions, adopting a topological viewpoint holds significant potential for unifying the behaviors at rational and irrational points, offering deeper insights into the structure of the mapping and streamlining practical computations.
@@ 2. Preliminaries of continued fractions
In this section, we list some basic properties of continued fractions, most of which can be found in the classical textbook on number theory by Hardy and Wright [4], as well as in books more focused on the theory of continued fractions by Iosifescu and Kraaikamp [5] and Khinchin [6].
For each $x\in[0,1]$, define
@@
Define $f\colon[0,1]\to\mathbb{N}_{\infty}^{\mathbb{N}}$ by
@@
for each $x\in[0,1]$. Clearly, $f$ is well-defined by definition. Moreover, $f$ extends the continued fraction mapping, as it satisfies $f(x)=(d_{k}(x))_{k\in\mathbb{N}}$ for any irrational $x\in(0,1)$. For a rational $x\in(0,1)$ with the continued fraction expansion $x=[d_{1}(x),\dotsc,d_{n}(x)]$, $f(x)$ is given by $(d_{1}(x),\dotsc,d_{n}(x),\infty,\infty,\infty,\dotsc)$ with $d_{n}(x)\geq 2$. If we adopt the convention $1/\infty\coloneqq 0$, it is evident that $f$ consistently yields the partial quotients.
@@
$\widetilde{\varphi}_{k}(\sigma)\coloneqq[\underbrace{\sigma_{1},\sigma_{2},\dotsc,\sigma_{k}}_{\text{$k$ terms}},\infty,\infty,\dotsc]=\cfrac{1}{\sigma_{1}+\cfrac{1}{\sigma_{2}+\cfrac{1}{\ddots+\cfrac{1}{\sigma_{k}}}}}$
for each $k\in\mathbb{N}_{0}$. We emphasize that $\widetilde{\varphi}_{k}$ is defined on the entire space $\mathbb{N}_{\infty}^{\mathbb{N}}$ rather than being restricted to $\Sigma_{k}$. For example, $\widetilde{\varphi}_{2}(3,\infty,4,\infty,\infty,\dotsc)=[3,\infty,\infty,\dotsc]=1/3$, and $\widetilde{\varphi}_{2}(\infty,\infty,\infty,\dotsc)=[\infty,\infty,\dotsc]=0$. Now, define $\widetilde{\varphi}\colon\mathbb{N}_{\infty}^{\mathbb{N}}\to[0,1]$ by
@@
$\varphi_{n}(\sigma)\coloneqq p_{n}^{*}(\sigma)/q_{n}^{*}(\sigma),$
@@
We remark that, in the traditional sense, every rational number in $(0,1)$ admits two distinct continued fraction expansions (see [4, Theorem 162] or [5, p. 3]). This is because, for any integer $k\geq 2$, we have $1/k=1/[(k-1)+1/1]$ with $k-1\in\mathbb{N}$. The following proposition rephrases this fact in terms of two maps $f\colon[0,1]\to\Sigma$ and $\varphi\colon\Sigma\to[0,1]$.
@@
    $\sigma$ $\coloneqq(\underbrace{\overline{d}_{1}(x),\overline{d}_{2}(x),\dotsc,\overline{d}_{n-1}(x)}_{\text{$n-1$ terms}},\overline{d}_{n}(x),\infty,\infty,\dotsc)=f(x)\in\Sigma_{n}\cap f([0,1]),$
    $\sigma^{\prime}$ $\coloneqq(\underbrace{\overline{d}_{1}(x),\overline{d}_{2}(x),\dotsc,\overline{d}_{n-1}(x)}_{\text{$n-1$ terms}},\overline{d}_{n}(x)-1,1,\infty,\infty,\dotsc)\in\Sigma_{n+1}\setminus f([0,1]),$
@@
We define the fundamental interval of length $n$ for all $n\geq 1$ and $\sigma\in\Sigma$ by
[…]
@@ 3. Main results and proofs
For the first assertion, we note that $\sum_{k=1}^{\infty}1/F_{k}^{2}<\infty$, which follows directly from Binet’s formula (1.1). In fact, since $\phi\coloneqq(1+\sqrt{5})/2>1$ and $(1-\sqrt{5})/2=-1/\phi$ has absolute value less than $1$, we have the convergence $(1/F_{k}^{2})/(1/\phi^{2k})\to 5$ as $k\to\infty$. Moreover, $\sum_{k=1}^{\infty}(1/\phi^{2k})<\infty$ because it is a geometric series with ratio $11/\phi^{2}<$ 1. By the limit comparison test, it follows that $\sum_{k=1}^{\infty}1/F_{k}^{2}<\infty$. We omit the remaining details, as the proof follows the same line as the standard proof that any countable product of metric spaces is metrizable. The second statement follows from Tychonoff’s theorem, which tells us that the product space $\mathbb{N}_{\infty}^{\mathbb{N}}$ is compact. ∎
@@
Note first that, for any $\sigma\in\mathbb{N}_{\infty}^{\mathbb{N}}$, we have $g(\sigma)\in\Sigma$, and so $\varphi(g(\sigma))=\widetilde{\varphi}(g(\sigma))$ by definition of $\varphi$ in (2.1). If $\sigma\in\Sigma_{\infty}$, then $g(\sigma)=\sigma$ by definition, and hence, $\widetilde{\varphi}(g(\sigma))=\widetilde{\varphi}(\sigma)$. Now, assume $\sigma\in\mathbb{N}_{\infty}^{\mathbb{N}}\setminus\Sigma_{\infty}$, and put $\sigma\coloneqq(\sigma_{k})_{k\in\mathbb{N}}$. Then, we can find the unique $n\in\mathbb{N}_{0}$ such that $\sigma^{(n)}\in\Sigma_{n}$ and $\sigma_{n+1}=\infty$, so that $g(\sigma)=\sigma^{(n)}\in\Sigma_{n}$. It follows by definition that $\widetilde{\varphi}(\sigma)=\widetilde{\varphi}_{n}(\sigma)=\widetilde{\varphi}_{n}(g(\sigma))=\widetilde{\varphi}(g(\sigma))$, and this completes the proof. ∎
@@
Let $\sim_{g}$ be the binary relation on $\mathbb{N}_{\infty}^{\mathbb{N}}$ given by $\sigma\sim_{g}\tau$ if and only if $g(\sigma)=g(\tau)$ for $\sigma,\tau\in\mathbb{N}_{\infty}^{\mathbb{N}}$. Evidently, $\sim_{g}$ is an equivalence relation on $\mathbb{N}_{\infty}^{\mathbb{N}}$. Note that the map $\widetilde{g}\colon\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}\to\Sigma$ given by $\widetilde{g}([\sigma])=g(\sigma)$, for each $[\sigma]\in\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}$, is a bijection. We equip $\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}$ with the quotient topology, and define
@@
Case I. Assume $\tau_{M}\neq\infty$. Since the first $M$ terms of $g(\sigma)$ are all finite, we have $\sigma^{(M)}=(g(\sigma))^{(M)}\in\Sigma_{M}$ by definition of $g$; similarly, $\tau^{(M)}=(g(\tau))^{(M)}\in\Sigma_{M}$. Note that $\sigma_{M}\neq\tau_{M}$ by definition of $M$. Since $\bm{\sigma}_{k}\to\sigma$ and $\bm{\tau}_{k}\to\tau$ as $k\to\infty$ in the product space $\mathbb{N}_{\infty}^{\mathbb{N}}$, we know that $(\bm{\sigma}_{k})^{(M)}=\sigma^{(M)}\in\Sigma_{M}$ and $(\bm{\tau}_{k})^{(M)}=\tau^{(M)}\in\Sigma_{M}$ for all large enough $k$. Then, for such $k$, it must be that $(g(\bm{\sigma}_{k}))_{M}=\sigma_{M}$ and $(g(\bm{\tau}_{k}))_{M}=\tau_{M}$. But $\sigma_{M}\neq\tau_{M}$, and it follows that $g(\bm{\sigma}_{k})\neq g(\bm{\tau}_{k})$, i.e., $\bm{\sigma}_{k}\not\sim_{g}\bm{\tau}_{k}$, for all large enough $k$, a contradiction.
@@
Let $X$ be a compact space and $\sim$ an equivalence relation on $X$. If $X/\sim$ denotes the quotient space, then the following statements are equivalent:
@@
We first note that, by Lemma 3.7 and Proposition 3.8, the canonical projection $\pi_{g}\colon\mathbb{N}_{\infty}^{\mathbb{N}}\to\mathbb{N}_{\infty}^{\mathbb{N}}/\sim_{g}$ is a closed mapping. But $g=\widetilde{g}\circ\pi_{g}$ by Lemma 3.6, where $\widetilde{g}$ is a homeomorphism (Lemma 3.5), and this implies that $g$ is also a closed mapping.
@@
Theorem 3.12.
Proof.
$\iff\varphi^{-1}(U)\text{ is open in }(\Sigma,\mathcal{T}_{\Sigma}),$
@@
    $\lim_{\begin{subarray}{c}t\to x\\ t\in I_{n}(\sigma)\end{subarray}}f(t)=\sigma\quad\text{and}\quad\lim_{\begin{subarray}{c}t\to x\\ t\not\in I_{n}(\sigma)\end{subarray}}f(t)=\tau,$
    where $\sigma\in\Sigma_{n}$ for some $n\in\mathbb{N}$.
@@
(i) In view of Proposition 2.4(i), it is sufficient to show that $f$ is continuous at points $x\in[0,1]$ where $\varphi^{-1}(\{x\})$ is a singleton. Suppose, for contradiction, that $f$ is not continuous at such a point $x$. Then, there exists a sequence $(x_{k})_{k\in\mathbb{N}}$ converging to $x$ such that the corresponding images $f(x_{k})$ fail to converge to $f(x)$. Using the compact metrizability of $\Sigma$ (Lemma 3.10), we can extract a subsequence of $(f(x_{k}))_{k\in\mathbb{N}}$ that converges to some $\tau\in\Sigma$. The continuity of $\varphi$ (Theorem 3.12) implies that $x=\varphi(\tau)$, and the singleton assumption forces $\tau=f(x)$. This contradicts the initial choice of the sequence $(x_{k})_{k\in\mathbb{N}}$, establishing the continuity of $f$ at such points.
(ii) Suppose, to the contrary, that $\lim_{\begin{subarray}{c}t\to x\\ t\in I_{n}(\sigma)\end{subarray}}f(t)\neq\sigma$. Then, there exists a sequence $(x_{k})_{k\in\mathbb{N}}$ in $I_{n}(\sigma)$ converging to $x$ such that $(f(x_{k}))_{k\in\mathbb{N}}$ does not converge to $\sigma$. By the compact metrizability of $\Upsilon_{\sigma}$ (Lemma 3.11), we extract a subsequence of $(f(x_{k}))_{k\in\mathbb{N}}$ converging to some $\upsilon\in\Upsilon_{\sigma}$. The continuity of $\varphi$ (Theorem 3.12) implies that $x=\varphi(\upsilon)$. The doubleton assumption forces $\upsilon=\sigma$ or $\upsilon=\tau$. Since $\tau\not\in\Upsilon_{\sigma}$, it must be that $\upsilon=\sigma$, which contradicts the initial assumption about the sequence $(x_{k})_{k\in\mathbb{N}}$. Thus, $\lim_{\begin{subarray}{c}t\to x\\ t\in I_{n}(\sigma)\end{subarray}}f(t)=\sigma$.
The proof for the second equality is similar. One needs to consider sequences in $\Sigma\setminus\Upsilon_{\sigma}$ and uses the compact metrizability of this subspace (Lemma 3.11). ∎
@@
It is a standard fact that the sets $\mathbb{I}$ and $\Sigma_{\infty}$ are in a bijective relation via the continued fraction mapping. Therefore, to obtain the corollary, we only need to combine Theorems 3.13(i) and 3.12. ∎
In conclusion, the results presented in this paper provide a topological framework that unifies the continuity behavior of the continued fraction mapping at both irrational and rational points. As mentioned in the introduction, this framework can be leveraged to reestablish the results of [7], specifically regarding the continuity of the error-sum function at irrational points and the one-sided continuity of the function at rational points, without relying on the complicated computations traditionally required. While the detailed derivation of these results using the topological approach is beyond the scope of this paper, readers are encouraged to explore this connection further to appreciate the utility of the framework introduced here.
Acknowledgements
I would like to express sincere gratitude to the referees for their thorough and insightful review of this paper. Their detailed comments and suggestions have immensely improved the manuscript, particularly in terms of clarity, rigor, and readability.